One card is selected randomly from 26 cards bearing the English alphabet, one distinct letter per card. What is the probability that the letter is E or S?
There are 26 equally likely outcomes. Exactly two distinct outcomes, E and S, satisfy the condition; the events are disjoint. Their probabilities add to $1/26+1/26=2/26$. Therefore **A**.
More Probability and Statistics questions
GATE 2026 ES Q15– Which expression is Pearson’s chi-square statistic? Here O i,E i are observed and expected…
GATE 2026 ES Q48– PM concentrations at seven locations are 45,40,55,65,60,35,50 µg/m³. Find their standard…
GATE 2025 ES Q24– Consider (i) Pollutant concentrations are generally modeled as lognormal;(ii)…
GATE 2024 ES Q11– Ten cards are numbered1–10. Find the probability of drawing an even number or a multiple…
GATE 2024 ES Q36– Ten observations are drawn from a normal population with mean µ. Their mean is x and…
GATE 2023 ES Q30– Disease prevalence is .002. A test is positive with probability.98 given disease and .03…